# Jyler Menard,a and RA Mansbacha

### v2

@@
[Uncaptioned image] Low-dimensional semi-supervised latent Bayesian optimization for designing antimicrobial peptides
@@
[Uncaptioned image] Antimicrobial peptides (AMPs) are a promising class of therapeutics to treat bacterial infections. Discovering and designing such peptides is difficult because of the vast number of possible sequences of amino acids. Deep generative models, such as variational autoencoders, have shown value in peptide design due to their ability to model sequence space with a continuous-valued latent space. Although such models have already been used to great effect in biomolecular design, they still suffer from a lack of interpretability and rigorous quantification of latent space quality as a search space. We investigate (1) whether searching through a dimensionally-reduced variant of the latent design space may facilitate optimization, (2) how organizing latent spaces with physicochemical properties may improve the efficiency of optimizing antimicrobial activity, and (3) the interpretability of the spaces. We find that employing a dimensionally-reduced version of the latent space is more interpretable and can be advantageous, while we can organize the latent space with different physicochemical properties even at different percentages of available labels. This work lays crucial groundwork for biophysically-motivated peptide design procedures.
footnotetext: a Department of Physics, Concordia University, Montréal, QC, H4B 1R6, Canada.    E-mail: re.mansbach@concordia.cafootnotetext: † Electronic Supplementary Information (ESI) available: [details of any supplementary information available should be included here]. See DOI: 00.0000/$0000$.
@@ 1 Introduction
Antimicrobial resistance is a growing global health concern. Through increasing usage in animal agriculture 1, aquaculture 2, and in humans 3, it is expected that by 2050, >8.2>8.2 million deaths per year will be associated with AMR 4. With traditional pharmaceutical companies becoming less productive 5, fewer antibiotics being developed 6, 7, and of those that have been developed since 2017, 10/13 target mechanisms with known resistance 8, there is an urgent need for novel classes of therapeutics to treat bacterial infections, and methods to quickly find them.
@@
One potentially fruitful area of research is the design and development of new and more effective antimicrobial peptides (AMPs). Antimicrobial peptides are short proteins of around 5 to 100 residues with broad-spectrum antibacterial activity 9. Many AMPs are thought to act by disrupting the bacterial membrane, through the formation of large pores, through inducing micellization of membrane patches (the so-called “carpet method”), or through a combination of such mechanisms 10. Because of the comparatively indiscriminate nature of the entire membrane as a target rather than a small protein binding pocket like most traditional small-molecule drugs, it was hypothesized that resistance to membrane-active AMPs will develop more slowly than to many traditional small-molecule antibiotics. Evidence supporting this hypothesis includes in vitro experiments across four different gram-negative bacterial species showing less resistance developing to membrane-active antibiotics than to traditional antibiotics within the same time frame 11. Another study also demonstrated that strong resistance to traditional antibiotics appeared within 120 generations of repeated exposure, compared to the 600 generations required for resistance to develop to an analogue of the peptide Magainin-2 12. In addition to the possibility of slowing resistance, certain AMPs can be used as part of combination drug cocktails to improve the efficacy of other antibiotics 13. Such combination therapies have also been found to have slower resistance development14. Due to their broad-spectrum activity, slow resistance, synergism with other antibiotics, and the need for new antibiotics, AMPs are a promising therapeutic class.
@@
While peptides are short compared to other proteins, the space of all possible peptide sequences is still vast. Considering only naturally-occurring amino acids, there are on the order of $\sum_{L=5}^{100}20^{L}$ possible sequences. Traditional mutagenesis methods are inadequate to explore more than a tiny fraction of this sequence space to find peptides with properties of interest, while traditional computational methods, such as virtual screening, are restricted to currently available protein datasets and cannot suggest completely novel candidate AMPs. Because of the number of possible peptides compared to the speed at which they can be tested, finding peptides with desired properties is difficult.
@@
Deep learning generative models can partially obviate the need to pre-generate libraries of sequences by providing a method to sample arbitrarily many new sequences. Recently, such models have been adopted by the biomolecular design community and have demonstrated a powerful capacity to produce peptides with desirable properties15, 16, 17, 18, 19, 20. Thus far, however, the focus of these models has tended to be on design of peptides for specific applications rather than an understanding of the ways in which the specific model and procedure affect the underlying design space. Recently, we developed quantitative benchmarks and investigated the properties of the design space itself in the context of our own generative model for antimicrobial peptide sequence design based on a variational autoencoder (VAE) 21. VAE-based models possess associated latent spaces that can be thought of as explicit continuous design space representations of the discrete peptide sequences. This formulation of the problem facilitates a focus on the quality of the design space itself. In this article, we expand on this work to investigate how the organization of the design space can affect optimization procedures in that space.
@@
One can employ BayesOpt to search through the latent space of an associated generative model; this is referred to as Latent Bayesian Optimization (LBO) 22, 23, 24, 25. LBO has been previously successfully applied to the design of small molecules 22 and synthetic optoelectronic peptides 26. Although screening peptides sampled from a generative model to design AMPs without using Bayesian optimization also has precedent 18, 19, we focus on LBO because it guides how to explore the design space and is a natural technique to employ if one is interested in evaluating the design spaces themselves.
@@
Other works have explored how the organization of the latent space can impact the efficiency of LBO. A number of ways to organize the latent space have been investigated, with particular focus on small molecule discovery where large amounts of data are readily available. In Gómez-Bombarelli et al. 22, the authors investigated whether jointly training a property predictor and a decoder can improve the efficiency of LBO for the task of finding optimal small molecules, finding compelling evidence that it can. Lee et al. 24 examined how LBO performs in a latent space induced to be correlated with the objective they are trying to optimize, finding promising results. In Grosnit et al. 23 the authors compare a number of contrastive learning methods, finding that a triplet loss can be effective. Given these works, it seems the efficiency of LBO can be improved by organizing the underlying latent space being searched through, but this has not been previously investigated in the context of antimicrobial peptide design.
@@
Unlike in the small molecule design space, where large amounts of data relevant to the performance of such molecules as antibiotics are readily available, the corresponding data on antimicrobial peptides is relatively sparse. There are hundreds of thousands or even millions of examples of small-molecule data as experimental drug-target binding affinities and well-established computed properties related to drug-likeness and solubility readily available. Conversely, in the context of peptides, antimicrobial activity in the form of experimentally-measured minimum inhibitory concentration (MIC) values is scarce; specifying particular bacterial species as targets restricts the available data even further. For example, for E. coli there are just 11k entries with corresponding MIC measurements, and if we restrict ourselves to those sequences with more than one reported MIC measurement, then we are left with 4k entries 27. Experimentally-measured peptide hemolysis, which is important for the assessment of host toxicity, is available for even fewer examples, with DBAASP containing 8k sequences with reported hemolysis measurements that falls to 1.6k when restricting ourselves to more than one measurement per sequence 27. While this may be a sufficient amount for traditional machine learning methods, it is unclear how this amount may impact the effectiveness of generative deep learning models and the organization of their latent space. Moreover, it is unclear whether using easily-computed physicochemical properties as proxies may offer a benefit over the more direct but limited amount of activity data.
@@
In this work, we explore LBO for antimicrobial peptide design spaces. We introduce and assess a modification of LBO as it has been previously applied in drug discovery applications by performing Bayesian Optimization over a reduced space representation of the latent space for AMP design. Rather than performing Bayesian Optimization over the high-dimensional latent space representation, we explore the possibility of performing Bayesian Optimization over a much lower-dimensional projection of that space constructed with Principal Components Analysis. The purpose of this approach is twofold: (i) we expect an increase in efficiency by using BayesOpt in a lower-dimensional space, and (ii) we expect an increase in interpretability by performing optimization in a more similar space to that which we employ for visualization compared to using reduced projections solely to visualize (a proxy of) the design space.
@@
We further investigate three important questions related to optimizing measures of AMP efficacy. We first ask whether organizing the latent space with easily-computable physicochemical properties that are not directly related to AMP activity is more informative or higher-performing than organizing the latent space with sparse amount of AMP activity data. We secondly ask whether organizing by more properties leads to a more structured, easier to optimize in, latent space. We thirdly ask whether increasing the percentage of labelled data available has a direct relationship to LBO efficiency and latent space organization, or if there are diminishing returns. By providing an analysis of LBO in both a semi-supervised and latent space organization context, we hope to further the discussions of how best to perform Bayesian optimization in latent spaces in a manner that informs us about the design space that is being traversed by the optimization algorithm, and introduce the technique to the therapeutic peptide design problem.
@@ 2.1 Bayesian Optimization
To find highly active antimicrobial peptides, we use Bayesian Optimization ( BayesOpt ) with a single objective function, as implemented in the BoTorch Python package using the SingleTaskGP class 28. BayesOpt is an iterative method of efficiently identifying design optima by using a surrogate (usually a quicker, less accurate assessment) of a desired objective function in combination with an acquisition function that selects the next potential design to be evaluated according to the true objective function (usually slower and more accurate assessment)29. A good acquisition function balances exploration of the design space with exploitation of observed data points29; in other words, it balances assessing new regions in design space with searching near known high-performing designs.
@@
$M_{l}(x)$ $=\Sigma_{l}(\vec{x},x_{D})\Sigma_{l}(x_{D},x_{D})^{-1}\cdot(\vec{f}(x_{D})-{m_{l}}(x_{D}))+m_{l}(\vec{x})$ (4)
$\Sigma_{l}(x)$ $=\sigma_{l}(x,x)-\sigma_{l}(x,x_{D})\sigma_{l}(x_{D},x_{D})^{-1}\sigma_{l}(x_{D},x),$ (5)
where $\vec{f}$ is the true objective function, $x_{D}$ is the dataset of $n$ previously-assessed input points, $\Sigma_{l}(x_{D},x_{D})$ is an $n\times n$ matrix $\Sigma_{l}(x_{D},x_{D})=\big(k(\vec{x}_{i},\vec{x}_{j})\big)_{1\leq i\leq n,\;1\leq j\leq n}$, and $\Sigma_{l}(\vec{x},x_{D})$ is a vector of length $n$ $\Sigma_{l}(\vec{x},x_{D})=\big(k(\vec{x},x_{i})\big)_{1\leq i\leq n}$. We use the radial basis function kernel (default when using BoTorch’s SingleTaskGP 28), but other choices exist. Because the Radial Basis Function Kernel is distance-dependent, there is an implicit assumption that points near each other have similar objective values. Through the distance-dependence, points near each other in the design space will be predicted to have similar surrogate objective values. With the posterior distribution in hand, predictions are made either by sampling from the posterior distribution
$\hat{f}(x)\sim\mathcal{G}\big(M_{l}(x),\;\Sigma_{l}(x)\big)$ (6)
@@
$\text{EI}(x)\equiv\mathbb{E}_{\hat{f}(x)}\left[\left(\hat{f}(x)-\vec{f}_{l}^{*}\right)^{+}\right],$
where $\vec{f}_{l}^{*}$ is the best value we have observed up to iteration $l$, $\hat{f}(x)$ is the predicted value of the surrogate function at point $x$, and $\left(\hat{f}(x)-\vec{f}_{l}^{*}\right)^{+}=\max\left(0,\hat{f}(x)-\vec{f}_{l}^{*}\right)$ is the improvement. $\text{EI}(x)$ encourages exploitation by searching for a point that is likely to improve on the current best point, and it encourages exploration through the standard deviation and cumulative distribution of the GPR function.
For our case, rather than using the Expected Improvement, we use the Log Expected Improvement, which has been found to be nearly identical but is easier to numerically optimize30. It is calculated as the logarithm of the expectation value of the improvement over the previous best value according to the surrogate model,
$\text{LogEI}(x)=\log{\mathbb{E}_{\hat{f}(x)}[(\hat{f}(x)-\vec{f}_{l}^{*})^{+}]}$
@@
$x_{l+1}=\text{argmax}_{x}\left(\;\text{LogEI}(x)\;\right)$
After assessing the $x_{l+1}$-th point using the true objective function, we add the pair $(x_{l+1},\;\vec{f}(x_{l+1}))$ to the dataset $\mathcal{D}=\mathcal{D}\;\cup\{(x_{l+1},f(x_{l+1}))\}$ and start the next iteration.

### v3

@@
[Uncaptioned image] Towards best practices in low-dimensional semi-supervised latent Bayesian optimization for the design of antimicrobial peptides
@@
[Uncaptioned image] Generative deep learning techniques have demonstrated an impressive capacity for tackling biomolecular design problems in recent years. Despite their high performance, however, they still suffer from a lack of interpretability and rigorous quantification of associated search spaces, which are necessary to unlock their full potential for scientific inquiry beyond efficient design. An area in which they are of particular interest is in the design of antimicrobial peptides, which are a promising class of therapeutics to treat bacterial infections. Discovering and designing such peptides is difficult because of the vast number of possible sequences and comparatively small amount of experimental information. In this work, we perform a theoretical investigation of latent Bayesian optimization for searching through peptide sequence spaces, with a focus on antimicrobial peptides. We investigate (1) whether searching through a dimensionally-reduced variant of the latent design space may facilitate optimization, (2) how organizing latent spaces by differing amounts of more and less relevant information may improve the efficiency of arriving at an optimal peptide design, and (3) the interpretability of the spaces. We find that employing a dimensionally-reduced version of the latent space is more interpretable and can be advantageous, while the use of less-relevant but more easily-computable physicochemical properties is advantageous to latent space organization in certain contexts and the use of more-relevant but sparser properties associated with the latent Bayesian objective function is advantageous in others. This work lays crucial groundwork for biophysically-motivated peptide design procedures, with an especial focus on antimicrobial peptides.
footnotetext: a Department of Physics, Concordia University, Montréal, QC, H4B 1R6, Canada.    E-mail: re.mansbach@concordia.cafootnotetext: † Electronic Supplementary Information (ESI) available: Further detailed analyses, tables and figures available in PDF form (4 sections, 10 tables, and 24 figures). Code for model training and analysis can be found at https://github.com/Mansbach-Lab/compare-latent-spaces-amps/tree/main. Trained model checkpoints and datasets are available at https://doi.org/10.5281/zenodo.17872449 See DOI: 00.0000/$0000$.
@@ 1 Introduction
Antimicrobial resistance is a growing global health concern. Through increasing usage in animal agriculture 1, aquaculture 2, and in humans 3, it is expected that by 2050, >8.2>8.2 million deaths per year will be associated with AMR 4. With traditional pharmaceutical companies becoming less productive 5, fewer antibiotics being developed 6, 7, and of the thirteen that have been approved since 2017, ten belong to antibiotic groups with known resistance mechanisms 8, there is an urgent need for novel classes of therapeutics to treat bacterial infections, and methods to quickly find them.
@@
One potentially fruitful area of research is the design and development of new and more effective antimicrobial peptides (AMPs). Antimicrobial peptides are short proteins of around 5 to 100 residues with broad-spectrum antibacterial activity 9. Many AMPs are thought to act by disrupting the bacterial membrane, through the formation of large pores, through inducing micellization of membrane patches (the so-called “carpet method”), or through a combination of such mechanisms 10. Because of the comparatively indiscriminate nature of the entire membrane as a target rather than a small protein binding pocket like most traditional small-molecule drugs, it was hypothesized that resistance to membrane-active AMPs will develop more slowly than to many traditional small-molecule antibiotics. Evidence supporting this hypothesis includes in vitro experiments across four different gram-negative bacterial species showing less resistance developing to membrane-active antibiotics than to traditional antibiotics within the same time frame 11. In the same study, it was demonstrated that strong resistance to traditional antibiotics appeared within 120 generations of repeated exposure 11, compared to the 600 generations required for resistance to develop to an analogue of the peptide Magainin-2 that was observed by Perron et al. 12. Despite differences in study methodology, the evidence overall supports slower resistance development to peptides. In addition to the possibility of slowing resistance, certain AMPs can be used as part of combination drug cocktails to improve the efficacy of other antibiotics 13. Such combination therapies have also been found to have slower resistance development14. Due to their broad-spectrum activity, slow resistance, synergism with other antibiotics, and the need for new antibiotics, AMPs are a promising therapeutic class.
@@
While peptides are short compared to other proteins, the space of all possible peptide sequences is still vast. Considering only naturally-occurring amino acids, there are on the order of $\sum_{L=5}^{100}20^{L}$ possible sequences. Traditional mutagenesis methods are inadequate to explore more than a tiny fraction of this sequence space to find peptides with properties of interest. Traditional computational methods, including machine learning, can improve the speed of candidate identification by rapidly screening databases full of peptide and protein sequences. For example, some of the earliest predictive models for AMP identification were hidden markov models and shallow neural networks 15, 16, with logistic regression 17 and SVMs 18 following after. More recently, motivated by the success of deep neural networks for (natural language) sequence representations, deep neural network architectures began to be employed 19. All of these methods, however, function to identify candidates in a database and cannot suggest completely novel AMPs. Because of the number of possible peptides compared to the speed at which they can be tested, finding peptides with desired properties is difficult.
Deep learning generative models can partially obviate the need to pre-generate libraries of sequences by providing a method to sample arbitrarily many new sequences. Recently, such models have been adopted by the biomolecular design community and have demonstrated a powerful capacity to produce peptides with desirable properties. Capecchi et al. 20 trained a strain-specific RNN to generate peptide sequences and filtered them using classifiers. Li et al. 21 trained a generative model based on a transformer architecture intended to generate high-activity mutants from an input sequence. Zakharova et al. 22 used an RNN to generate anticancer peptides. Szymczak et al. 23 designed a conditional VAE–conditioned on categorical classifications of AMPness and high/low activity levels–in an attempt to directly sample from the high-activity AMP conditional probability distribution. Das et al. 24 used a Wasserstein autoencoder together with a number of classifiers to generate peptides.
@@
While impressive work is being done in improving both predictive and generative models, tying them together in a closed-loop optimization scheme is (currently) less common. Furthermore, the focus of these models has tended to be on design of peptides for specific applications rather than an understanding of the ways in which the specific model and procedure affect the underlying design space. Recently, we developed quantitative benchmarks and investigated the properties of the design space itself in the context of our own generative model for antimicrobial peptide sequence design based on a variational autoencoder (VAE) 25. VAE-based models possess associated latent spaces that can be thought of as explicit continuous design space representations of the discrete peptide sequences. This formulation of the problem facilitates a focus on the quality of the design space itself. In this article, we expand on this work to investigate how the organization of the design space can affect optimization procedures in that space.
@@
One can employ BayesOpt to search through the latent space of an associated generative model; this is referred to as Latent Bayesian Optimization (LBO) 26, 27, 28, 29. LBO has been previously successfully applied to the design of small molecules 26 and synthetic optoelectronic peptides 30. Although screening peptides sampled from a generative model to design AMPs without using Bayesian optimization also has precedent 23, 24, we focus on LBO because of its data efficiency and potential for application with higher fidelity, more costly oracles.
@@
Other works have explored how the organization of the latent space can impact the efficiency of LBO. A number of ways to organize the latent space have been investigated, with particular focus on small molecule discovery where large amounts of data are readily available. In Gómez-Bombarelli et al. 26, the authors investigated whether jointly training a property predictor and a decoder can improve the efficiency of LBO for the task of finding optimal small molecules, finding compelling evidence that it can. Lee et al. 28 examined how LBO performs in a latent space induced to be correlated with the objective they are trying to optimize, finding promising results. In Grosnit et al. 27 the authors compare a number of contrastive learning methods, finding that a triplet loss can be effective. Given these works, it seems the efficiency of LBO can be improved by organizing the underlying latent space being searched through, but this has not been previously investigated in the context of antimicrobial peptide design. One further consideration for LBO is that often the associated neural network has a relatively high-dimensional latent space. BayesOpt can have difficulty in high-dimensional spaces. On the other hand, reducing the dimensionality of the latent space may corrupt the ability of the associated generative model to generate diverse, novel, and realistic sequences. This leaves us with a dilemma where we may expect either worsened BayesOpt performance due to a high-dimensional latent space, or worsened generative model performance due to too small of a latent space.
@@
Unlike in the small molecule design space, where large amounts of data relevant to the performance of such molecules as antibiotics are readily available, the corresponding data on antimicrobial peptides is relatively sparse. There are hundreds of thousands or even millions of examples of small-molecule data as experimental drug-target binding affinities and well-established computed properties related to drug-likeness and solubility readily available. Conversely, in the context of peptides, antimicrobial activity in the form of experimentally-measured minimum inhibitory concentration (MIC) values is scarce; specifying particular bacterial species as targets restricts the available data even further. For example, for E. coli there are just 11k entries with corresponding MIC measurements, and if we restrict ourselves to those sequences with more than one reported MIC measurement, then we are left with 4k entries 31. Experimentally-measured peptide hemolysis, which is important for the assessment of host toxicity, is available for even fewer examples, with DBAASP containing 8k sequences with reported hemolysis measurements that falls to 1.6k when restricting ourselves to more than one measurement per sequence 31. While this may be a sufficient amount for traditional machine learning methods, it is unclear how this amount may impact the effectiveness of generative deep learning models and the organization of their latent space. Moreover, it is unclear whether using easily-computed physicochemical properties as proxies may offer a benefit over the more direct but limited amount of activity data.
@@
In this work, we determine best practices for performing latent Bayesian optimization in the context of peptide sequence spaces, with a focus on the antimicrobial peptide context in particular. We introduce and assess a modification of latent Bayesian Optimization as it has been previously applied in drug discovery applications by performing Bayesian Optimization over a reduced space representation of a peptide sequence-based the latent space for AMP design. Rather than performing Bayesian Optimization over the high-dimensional latent space representation, we explore the possibility of performing Bayesian Optimization over a much lower-dimensional projection of that space constructed with Principal Components Analysis. The purpose of this approach is twofold: (i) we expect an increase in efficiency by using BayesOpt in a lower-dimensional space without modifying the size of the generative model’s latent space, and (ii) we expect an increase in interpretability by performing optimization in a more similar space to that which we employ for visualization compared to using reduced projections solely to visualize (a proxy of) the design space. We also note that we employ a proxy objective function (an “oracle”) related to predicted AMP activity for better control of algorithmic design prior to deployment on more realistic, expensive objectives, such as experimental measurements of peptide activity.
@@
We further investigate three important questions related to incorporation of prior knowledge to latent Bayesian optimization in peptide sequence-based design spaces. We first ask whether organizing the latent space with easily-computable physicochemical properties that are not directly related to predicted AMP activity is more informative or higher-performing than organizing the latent space with a sparse amount of predicted AMP activity data. We secondly ask whether organizing by more properties leads to a more structured, easier to optimize in, latent space. We thirdly ask whether increasing the percentage of labelled data available has a direct relationship to LBO efficiency and latent space organization, or if there are diminishing returns. By providing an analysis of LBO in both a semi-supervised and latent space organization context, we hope to further the discussions of how best to perform Bayesian optimization in latent spaces in a manner that informs us about the design space that is being traversed by the optimization algorithm, and introduce the technique to the therapeutic peptide design problem in the context of AMPs.
@@ 2 Methods
Fig. 1: Schematic overview of three approaches to Latent Bayesian Optimization investigated herein (differences highlighted). LBO begins assuming a pre-trained generative model with an encoder and a decoder module. Each method follows a similar procedure: encode the optimization dataset into the latent space of a transVAE model; fit a Gaussian Process Regressor; maximize the Expected Improvement to select the next candidate to score using the oracle; decode the candidate latent point to a amino acid sequence; score the sequence using the oracle; add the now-labelled sequence to the optimization dataset, and repeat NN times. The differences lie in whether (a) we use the full latent space representation directly in fitting the GPR, (b) we use a PCA projection of the latent space when fitting the GPR, or (c) whether a small neural network that projects the latent points into a smaller-dimensional space is fit simultaneously with the GPR. Following (b) for one BayesOpt iteration: encode optimization dataset into the 64-dimensional latent space of a transVAE; project those points into a five-dimensional space constructed using PCA; fit a GPR and select a candidate five-dimensional point; inverse project back to the 64-dimensional latent space; decode the candidate latent point to a candidate sequence; score the sequence using the oracle.
@@ 2.1 Bayesian Optimization
To find highly active antimicrobial peptides, we use Bayesian Optimization ( BayesOpt ) with a single objective function, as implemented in the BoTorch Python package using the SingleTaskGP class 32. BayesOpt is an iterative method of efficiently identifying design optima by using a surrogate (usually a quicker, less accurate assessment) of a desired objective function in combination with an acquisition function that selects the next potential design to be evaluated according to the true objective function (usually slower and more accurate assessment)33. A good acquisition function balances exploration of the design space with exploitation of observed data points33; in other words, it balances assessing new regions in design space with searching near known high-performing designs.
@@
$M_{l}(\vec{x})$ $=m_{l}(\vec{x})+\sigma_{l}(\vec{x},x_{D})^{T}\sigma_{l}(x_{D},x_{D})^{-1}\cdot({f}(x_{D})-{m_{l}}(x_{D}))$ (4)
$\Sigma_{l}(\vec{x})$ $=\sigma_{l}(\vec{x},\vec{x})-\sigma_{l}(\vec{x},x_{D})^{T}\sigma_{l}(x_{D},x_{D})^{-1}\sigma_{l}(\vec{x},x_{D}),$ (5)
where f{f} is the true objective function, $x_{D}$ is short-hand for the dataset of $n$ previously-assessed input points $\{\vec{x}_{i}\}$, $\sigma_{l}(x_{D},x_{D})$ is an $n\times n$ matrix with entries $\big[\sigma_{l}(x_{D},x_{D})\big]_{i,j}=k(\vec{x}_{i},\vec{x}_{j})$, and $\sigma_{l}(\vec{x},x_{D})$ is a vector of length $n$ with entries $\big[\sigma_{l}(\vec{x},x_{D})\big]_{i}=k(\vec{x},\vec{x}_{i})$. We use the radial basis function kernel (default when using BoTorch’s SingleTaskGP 32), but other choices exist. Because the Radial Basis Function Kernel is distance-dependent, there is an implicit assumption that points near each other have similar objective values. Through the distance-dependence, points near each other in the design space will be predicted to have similar surrogate objective values. With the posterior distribution in hand, predictions are made either by sampling from the posterior distribution
$\hat{f}(\vec{x})\sim\mathcal{G}\big(M_{l}(\vec{x}),\;\Sigma_{l}(\vec{x})\big)$ (6)
@@
$\text{EI}(\vec{x})\equiv\mathbb{E}_{\hat{f}(\vec{x})}\left[\left(\hat{f}(\vec{x})-f_{l}^{*}\right)^{+}\right],$
where $f_{l}^{*}$ is the best value we have observed up to iteration $l$, $\hat{f}(x)$ is the predicted value of the surrogate function at point $\vec{x}$, and $\left(\hat{f}(\vec{x})-f_{l}^{*}\right)^{+}=\max\left(0,\hat{f}(\vec{x})-f_{l}^{*}\right)$ is the improvement. $\text{EI}(\vec{x})$ encourages exploitation by searching for a point that is likely to improve on the current best point, and it encourages exploration through the standard deviation and cumulative distribution of the GPR function.
For our case, rather than using the Expected Improvement, we use the Log Expected Improvement, which has been found to be nearly identical but is easier to numerically optimize34. It is calculated as the logarithm of the expectation value of the improvement over the previous best value according to the surrogate model,
$\text{LogEI}(x)=\log{\mathbb{E}_{\hat{f}(\vec{x})}[(\hat{f}(\vec{x})-f_{l}^{*})^{+}]}$
@@
$\vec{x}_{l+1}=\text{argmax}_{\vec{x}}\left(\;\text{LogEI}(\vec{x})\;\right)$
After assessing the $\vec{x}_{l+1}$-th point using the true objective function, we add the pair $(\vec{x}_{l+1},\;f(\vec{x}_{l+1}))$ to the dataset $\mathcal{D}=\mathcal{D}\;\cup\{(\vec{x}_{l+1},f(\vec{x}_{l+1}))\}$ and start the next iteration.